# Examining the Utility of Self-disclosure Types for Modeling Annotators of Social Norms

### v1

@@ Abstract
Recent work has explored the use of personal information in the form of persona sentences or self-disclosures to improve modeling of individual characteristics and prediction of annotator labels for subjective tasks. The volume of personal information has historically been restricted and thus little exploration has gone into understanding what kind of information is most informative for predicting annotator labels. In this work, we categorize self-disclosure sentences and use them to build annotator models for predicting judgments of social norms. We perform several ablations and analyses to examine the impact of the type of information on our ability to predict annotation patterns. We find that demographics are more impactful than attitudes, relationships, and experiences. Generally, theory-based approaches worked better than automatic clusters. Contrary to previous work, only a small number of related comments are needed. Lastly, having a more diverse sample of annotator self-disclosures leads to the best performance.
@@
Kieran Henderson  and Kian Omoomi  and Vasudha Varadarajan  and Allison Lahnala  and Charles Welch Toronto University, Carnegie Mellon University, McMaster University vvaradar@andrew.cmu.edu,{omoomik,lahnalaa,cwelch}@mcmaster.ca
@@ 1 Introduction
Our work addresses three research questions (RQs). RQ1. What strategies and parameters for sampling context about annotators produce the most effective models of their judgments of social situations? Specifically, we examine 1) how many samples are ideal for learning a person’s judgment style, 2) whether samples that are semantically similar to the social situations offer more predictive value than randomly selected ones, and 3) whether it is more effective to use selected sentences rather than entire posts. RQ2. How do different types of self-disclosures compare in how they influence the performance of annotator models? We used theory based regular expressions to categorize self-disclosure types into demographics (e.g., age, gender), experiences (e.g., hobbies, possessions, work life), attitudes (e.g., opinions), and relationships (e.g., family members, relationship status). RQ3. How do automatic categorizations of self-disclosure types compare to theory-based categorizations? We derive categorizations of self-disclosure types by a data-driven clustering method, and compare to our theory-based taxonomy.
@@ 2 Background
Particularly in subjective tasks, multiple valid perspectives may exist, making the use of a single ground truth problematic. When considering highly subjective tasks like offensive or toxic language detection, individual differences in demographics and personality significantly influence how people label dimensions of the data Leonardelli et al. (2021); Sang and Stanton (2022); Sap et al. (2022); Pei and Jurgens (2023); Frenda et al. (2023). Perspectivist approaches to these tasks not only help mitigate issues like silencing minority opinions in model representations, but also improve the predictive precision and robustness of models Cabitza et al. (2023), as was shown in research on hate speech detection Akhtar et al. (2020), and irony detection Frenda et al. (2023). Perspectivism in NLP, therefore, is a growing movement encouraging the construction and distribution of disaggregated datasets to embrace multiple, valid perspectives, and evaluation against individual labels rather than aggregated gold standards Frenda et al. (2024).
@@
One approach is to model individuals as annotators directly. Golazizian et al. (2024); Lee and Goldwasser (2022) introduced the Moral Foundations Subjective Corpus and built a multitask model that captures general task patterns and signals differences among initial annotators. They use a few-shot sample selection technique focused on capturing how new annotators’ behaviors differ from the general patterns, to learn individual perspectives. Mostafazadeh Davani et al. (2022) developed multi-annotator models, with varying degrees of shared parameters, finding that their multi-task model met or exceeded other methods on hate speech detection and emotion classification, as well as uncertainty estimates.
@@
Another approach is to group annotator characteristics and study perspective variance with respect to grouped variables. The significance of variance in subjective annotation tasks explained by demographic factors is demonstrated in an increasing number of works. For example, Pei and Jurgens (2023) identified variance by age, gender, race, and education levels of the annotators in question-answering, offensiveness, and politeness rating. Sap et al. (2022) measured dimensions of “attitudes” with respect to political and linguistic ideologies, showing variance in assessments of language toxicity rooted in these measures in addition to demographic factors including gender and race. However, Fleisig et al. (2024) pointed out that demographics are often used to explain some amount of disagreement, but other factors may help explain systematic disagreement. Deng et al. (2023) looked at eight datasets and found that only a portion of the variance in annotation can be explained by attributes including age, gender, geographic location, household income, political leaning, and education. Orlikowski et al. (2023) experimented with explicitly representing sociodemographic factors by adding group layers to the aforementioned multi-annotator models from Mostafazadeh Davani et al. (2022) to model individual annotator judgments; this did not significantly improve modeling performance, suggesting that variation may have more to do with individual differences.
@@
Recent work on automated detection of self-disclosures have leveraged taxonomies based on the sensitivity or ambiguity of the personal information revealed. For example, Bak et al. (2014) propose a self-disclosure topic model, categorizing conversational turns in social media as high (sensitive information), medium (non-sensitive information), or no self-disclosure. Barak and Gluck-Ofri (2007) define levels according to the degree of personal thoughts, feelings, and information in a statement. Valizadeh et al. (2021) built a classifier using levels for ambiguity or uncertainty of disclosure of symptoms in a medical setting.
@@ 3 Data
We use a set of posts from the Reddit community /r/amitheasshole (AITA). In the AITA forum, Reddit users post a description of a social situation and the actions they took and ask if the have done something wrong. Community members provide their judgments in the form of YTA (you’re the asshole) or NTA (not the asshole). We use the data from Plepi et al. (2024) which was derived from Welch et al. (2022)’s dataset of AITA situations (which in turn used a set of posts crawled by Forbes et al. (2020) for the Social-Chem-101 dataset) by filtering for authors that have between 20 and 500 additional comments somewhere else on the subreddit. Plepi et al. (2024)’s resulting dataset contains 21k posts written by 15k authors, and 212k verdicts provided by 13k annotators. These annotators also have an additional 3.6M comments from other parts of the subreddit excluding the 21k posts. The distribution is skewed toward around 70% NTA for all splits, as people prefer not to use the YTA label, or possibly because posters are biased in their portrayal of themselves. Each post has a title (situation description), full post text, author, and a set of comments from other Reddit users (commenters/annotators) which each provide a verdict (judgment YTA/NTA) and a justification or additional comment to supplement the verdict.
Our paper performs an annotator modeling task, wherein the aim is to predict how a user would “annotate” their verdict label given what we know about the author from self-disclosure statements. Collecting deeper self-disclosures as described by social penetration theory Tang and Wang (2012), is often expensive and challenging. While AITA posts are not formally "annotated" but rather "participated in" with some incentive, the platform’s anonymity and reduced real-world consequences make individuals more willing to share personal information De Choudhury and De (2014); Miller (2020). This makes AITA a practical proxy for large-scale annotation, offering a more affordable and relatively high-quality alternative.
@@
To investigate how different types of self-disclosures influence annotation modeling, we first sought to categorize annotator comments for the types of self-disclosures they contain. We attempted to assign both theory-based and automatic category assignments to each comment in our dataset to represent the type of information disclosed by annotators. A comment can belong to multiple theory-based categories, but can only belong to one cluster. Our theory-based assignment was informed in part by our intuition and related psychological theory, and in part by our exploratory analysis of the automatic assignments.
@@ 3.1 Automatic Categorization
We started with both LDA topics Blei et al. (2003) and k-means clusters as ways of grouping comments. When clustering and topic modeling, we use only posts that contain a self-disclosure from the list in Appendix A so that clusters relate more to personal statements. This leaves us with 570k posts, or 15.7% of the original data. About 400 authors have never written a post with a self-disclosure phrase and were excluded from our category experiments. Clusters were created using post embeddings generated using SBERT with the same parameters discussed in §4. Both approaches provide a method by which a single group label (topic or cluster) can be assigned to each comment. We found through manual inspection in preliminary experiments that both provided reasonable groupings and decided to proceed with k-means, as it is more easily interpretable. We also experimented with changing the number of clusters by multiples of five up to 25, finding that ten provided a reasonably interpretable and coherent set of clusters. We computed an average silhouette score of 0.25 for our resulting clusters Rousseeuw (1987). See Appendix D for more details.
@@
Table 4: Situation split performance (Accuracy and F1) across varying numbers of maximum samples, with best performance in bold. Similar 5 comments to random 5 comments statistically significant (t-test, $p<0.0002$).
@@ 3.2 Theory-based Categorization
Building upon the Social Penetration Theory for self disclosure Altman (1973), we introduce a theory-based categorization framework for understanding the types of self-disclosure that occur during interpersonal communication. Our taxonomy of annotator self-disclosure statements includes (a) Demographics, which was found to be the most common in other studies on annotator modeling, (b) Experiences, which includes generally self-descriptive statements about one’s past experiences and relatively more stable traits, (c) Attitudes, which are self-descriptive statements social views, morals, values and belief systems, and (d) Relationships, which is the sociability of the annotator. The clustering methods in all preliminary experiments picked up on the relationship between those in conflict, similar to what was found in Welch et al. (2022). We analyzed the most common n-grams around self-disclosure phrases, grouped authors by post count, tagged parts of speech to identify frequent patterns and verbs, and used Stanza’s coreference resolution (Qi et al., 2020) to clarify sentence meaning and better understand the main themes in self-disclosure.Theory-based and automatic approaches group self-disclosures by the type of information and the context (e.g. affect or conflict type) in which the disclosure was made, respectively. Post and author counts for each category are listed in Tables 2 and 3. We extended the age and gender regular expressions from Welch et al. (2020) and Plepi et al. (2022) to include a variety of ways of expressing age, non-binary and transgender, along with age and gender combinations, e.g. (24F representing 24 year old female). Our set of initial self-disclosure statements was expanded from Plepi et al. (2024) and similar to Mazaré et al. (2018), but without enforcing part-of-speech constraints. We also borrowed expressions from Tigunova et al. (2020), which included age, gender, family status, profession, and hobby, but expanded the set of cases. The full set of regular expressions we developed is listed in Table 6 in the Appendix.
@@
After using regex to extract the self-disclosure information, we randomly selected 100 comments from each group to annotate by hand. We found that there were 9 comments in the demographic group, 17 posts in the attitudes group, 25 in the experiences group, and 11 in the relationships group which were not clean self-disclosure phrases. For example, the attitudes regex patterns currently capture the phrases: "I think driving is scary" and "I think thats a good plan". The first phrase gives a clear opinion and is fully contained. The second sentence gives an opinion on something that has been mentioned outside of the current comment. We were initially expecting to extract clear opinions with this pattern, but in reality self-contained disclosures are difficult to extract.
@@ 4 Methodology
As a baseline, we reproduced the results of Plepi et al. (2022) for their situation split using the same parameters from their setup. They showed that the situation split was the most challenging, i.e. predicting how annotators respond to new conflict posts (see Appendix E for details on other splits). We used DistilRoBERTa Sanh et al. (2019) for the base SBERT classifier Reimers and Gurevych (2019) with a hidden size 768 and max length 512. We use an added linear layer for classification, trained the model for 10 epochs with the Adam optimizer, learning rate 1e31e-3, and focal loss Lee et al. (2021).
Experimental Setup. The brief exploration of self-disclosure statements in previous work on generation raised several questions for us. RQ1: Which sampling parameters perform best? First, how many samples are ideal for classifying judgments? Are similar samples more useful than random? Lastly, is selecting sentences more effective than taking the whole posts? We hypothesized that comments similar to the post we are classifying should perform better. Given the results of Plepi et al. (2024), we assumed that 15-20 similar posts would work best. However, we hypothesized that sentences may work better because they focus on only the important information and eliminate potential noise in longer posts.
@@
Our next question was whether or not manually categorizing personal information was necessary, or if we could rely on automatic methods. RQ3: Does automatically categorizing self-disclosures via clustering perform as well as theory-based categorization? We hypothesized that clusters would not perform as well as theory-based categories due to difficulty in enforcing that the similarity of clustered posts was primarily due to information they provided about their author.
To perform the first set of experiments we used all comments from all annotators, meaning that some randomly sampled or similar posts/sentences may not belong to any theory-based or automatic category. We chose to investigate the sampling parameters first, as the search space grows exponentially if we combine this with RQ2 and 3. For the latter research questions, we sampled only similar posts from one category at a time. For each category, we use the same set of annotators, however, each annotator does not have the same number of comments in each category, so we used however many were available up to five.
All of the models in our experiments were trained using a single NVIDIA H100 GPU. Generating similar comment embeddings and sentence tests took approximately 5 hours and only needed to be computed once each. Generating the embeddings for each category took less than a minute. A new model was needed for each of our experimental settings, a total of 24 models for the experiments in Table 4, a total of 14 models for the experiments in Table 5. Each model took approximately 2 hours to train and was averaged over 5 runs.
@@
Table 5: Results for both theory based and automatic categorization of posts on the situation split. 5+% denotes the percentage of annotators that had at least 5 posts for each category. All comments baseline represents previous work by Plepi et al. (2022). Demographics significantly outperforms both baselines (t-test, $p<0.0001$).
5 Results
The results of our first set of experiments are displayed in Table 4, showing accuracy and macro F1 score. We find that sentences underperform comments. After a certain extent, adding more sentences or comments stops adding new information. Text is sorted by similarity, thus there is a trade-off. When adding more similar text, we are adding less relevant information. When adding similar sentences, we are adding only the smallest unit of most similar information, yet adding 25-30 sentences already shows a drop in performance as each sentence is less similar and less useful than the last. Overall, we find that adding the 5 most similar comments performs best. Using similar information, not surprisingly, is more effective than adding random posts or sentences. We found that 5 similar comments significantly outperforms a baseline which uses all posts and which uses no posts (t-test, $p<0.0001$). We use the setting of 5 most similar comments for subsequent experiments.
Results for our experiments with categories are shown in Table 5, with highest accuracy and F1 in bold separately for theory-based and automatic categories. The highest performing theory-based category was demographics, followed by attitudes, and finally experiences, and relationships. We compared to a baseline that uses no comments from the annotators and a baseline that uses all comments, as in previous work. Notably, using all comments performs worse, as we are not filtering for useful information. The highest performing automatic clusters were C7: Sympathy & Support and C9: Negative Affect. C3: Parenting & Discipline had a notably high accuracy, but not significantly different F1 score, though C3 and C5: Relationship Issues with Men tied for the next highest scoring clusters. Notably, C7 had one of the lowest coverage rates with only 24% of authors having five or more posts in that cluster.

### v2

@@ Abstract
Recent work has explored the use of personal information in the form of persona sentences or self-disclosures to improve modeling of individual characteristics and prediction of annotator labels for subjective tasks. The volume of personal information has historically been restricted and thus little exploration has gone into understanding what kind of information is most informative for predicting annotator labels. In this work, we categorize self-disclosures and use them to build annotator models for predicting judgments of social norms. We perform several ablations and analyses to examine the impact of the type of information on our ability to predict annotation patterns. Contrary to previous work, only a small number of comments related to the original post are needed. Lastly, a more diverse sample of annotator self-disclosures did not lead to the best performance. Sampling from a larger pool of comments without filtering still yields the best performance, suggesting that there is still much to uncover in terms of what information about an annotator is most useful for verdict prediction.
@@
Kieran Henderson  and Kian Omoomi  and Vasudha Varadarajan  and Allison Lahnala  and Charles Welch University of Toronto, Carnegie Mellon University, McMaster University hende325@cs.toronto.edu,vvaradar@andrew.cmu.edu,{omoomik,lahnalaa,cwelch}@mcmaster.ca
@@ 1 Introduction
Our work addresses three research questions. RQ1. What strategies and parameters for sampling information about annotators produces the most effective models of an annotators judgments of social situations? Specifically, we examine 1) how many samples are ideal for learning a person’s judgment style, 2) whether samples that are semantically similar to the social situations offer more predictive value than randomly selected ones, and 3) whether it is more effective to use selected sentences rather than entire comments. RQ2. How do different types of self-disclosures compare in how they influence the performance of annotator models? We used theory based regular expressions to categorize self-disclosure types into demographics (e.g., age, gender), experiences (e.g., hobbies, possessions, work life), attitudes (e.g., opinions), and relationships (e.g., family members, relationship status). RQ3. How do automatic categorizations of self-disclosure types compare to theory-based categorizations? We derive categorizations of self-disclosure types by data-driven clustering methods, and compare to our theory-based taxonomy.
@@ 2 Background
Particularly in subjective tasks, multiple valid perspectives may exist, making the use of a single ground truth problematic. When considering highly subjective tasks like offensive or toxic language detection, individual differences in demographics and personality significantly influence how people label dimensions of the data Leonardelli et al. (2021); Sang and Stanton (2022); Sap et al. (2022); Pei and Jurgens (2023); Frenda et al. (2023). Perspectivist approaches to these tasks not only help mitigate issues like silencing minority opinions in model representations, but also improve the predictive precision and robustness of models Cabitza et al. (2023), as was shown in research on hate speech detection Akhtar et al. (2020), and irony detection Frenda et al. (2023). Even in “seemingly objective tasks” Basile et al. (2021); Frenda et al. (2024), like coreference resolution Recasens et al. (2011) or part-of-speech tagging Plank et al. (2014), nuanced models with a continuum/variety of perspectives may more effectively capture the linguistic phenomena. Perspectivism in NLP, therefore, is a growing movement encouraging the construction and distribution of disaggregated datasets to embrace multiple, valid perspectives, and evaluation against individual labels rather than aggregated gold standards Frenda et al. (2024).
@@
One approach to perspectivism is to learn from the disagreements themselves. Fornaciari et al. (2021) present one way to do this, by predicting a soft-label distribution, or the distribution of annotator responses, as an auxiliary task for part-of-speech tagging and stemming, finding that it leads to improved performance. One approach is to model individuals as annotators directly rather than collapsing their judgments into a single aggregated label. For example, Golazizian et al. (2024); Lee and Goldwasser (2022) introduced the Moral Foundations Subjective Corpus and proposed a multitask architecture that captures general task-level patterns and preserves differences among annotators. Their approach uses a few-shot sample selection technique to capture how new annotators’ labeling behaviors differ from the general patterns, which enables the learning of individual perspectives. Similarly, Mostafazadeh Davani et al. (2022) developed multi-annotator models with varying degrees of shared parameters, finding that explicitly accounting for annotator specific signals can match or exceeded aggregation based methods on subjective tasks such as hate speech detection and emotion classification, while also producing more informative uncertainty estimates.
@@
Another approach is to group annotators by shared characteristics and study perspective variance with respect to grouped variables. An increasing number of works demonstrate that a significant portion of disagreement in subjective annotation tasks can be explained by demographic factors. For example, Pei and Jurgens (2023) identified systematic variance associated with annotators’ age, gender, race, and education levels in tasks such as question-answering, offensiveness, and politeness rating. Similarly, Sap et al. (2022) operationalize annotator attitudes related to political and linguistic ideologies, showing that variation in language toxicity judgments is rooted in these ideological measures in addition to demographic factors such as gender and race. However, Fleisig et al. (2024) point out that demographics are often used to explain some amount of annotator disagreement, but other factors may better account for differences in perspectives. Along similar lines, Deng et al. (2023) analyzed eight datasets and found that only a limited portion of annotator variance can be explained by observable attributes such as age, gender, geographic location, household income, political leaning, and education. Orlikowski et al. (2023) explicitly incorporated sociodemographic factors into multi-annotator models by extending the architecture of Mostafazadeh Davani et al. (2022) with group level layers. Their results showed no significant performance gains, suggesting that perspective variation may be driven more by individual differences than by coarse demographic groupings.
@@
Recent work on automated detection of self-disclosures has leveraged taxonomies based on the sensitivity or ambiguity of the personal information revealed. For example, Bak et al. (2014) propose a self-disclosure topic model, categorizing conversational turns in social media as high (sensitive information), medium (non-sensitive information), or no self-disclosure. Barak and Gluck-Ofri (2007) define levels according to the degree of personal thoughts, feelings, and information in a statement. Valizadeh et al. (2021) built a classifier using levels for ambiguity or uncertainty of disclosure of symptoms in a medical setting.
@@ 3 Data
We use a set of posts from the Reddit community /r/amitheasshole (AITA). In the AITA forum, Reddit users post a description of a social situation and the actions they took and ask if they have done something wrong. Community members provide their judgments in the form of YTA (you’re the asshole) or NTA (not the asshole). We use the data from Plepi et al. (2024) which was derived from Welch et al. (2022)’s dataset of AITA situations (which in turn used a set of posts crawled by Forbes et al. (2020) for the Social-Chem-101 dataset) by filtering for authors that have between 20 and 500 additional comments somewhere else on the subreddit. Plepi et al. (2024)’s resulting dataset contains 21k posts written by 15k authors, and 212k verdicts provided by 13k annotators. These annotators also have an additional 3.6M comments from other parts of the subreddit excluding the 21k posts. The distribution is skewed toward NTA, with around 70% of all verdicts being NTA for all splits. This may be because people prefer not to use the YTA label, or possibly because posters are biased in their portrayal of themselves. Each post has a title (situation description), full post text, author, and a set of comments from other Reddit users (commenters/annotators) which each provide a verdict (judgment YTA/NTA) and a justification or additional comment to supplement the verdict.
Our paper performs an annotator modeling task, wherein the aim is to predict how a user would “annotate” their verdict label given what we know about the author from self-disclosure statements. Collecting deeper self-disclosures as described by social penetration theory Tang and Wang (2012), is often expensive and challenging. While AITA posts are not formally “annotated” but rather “participated in” with some incentive, the platform’s anonymity and reduced real-world consequences make individuals more willing to share personal information De Choudhury and De (2014); Miller (2020). This makes AITA a practical proxy for large-scale annotation, offering a more affordable and relatively high-quality alternative.
@@
To investigate how different types of self-disclosures influence annotation modeling, we first sought to categorize annotator comments for the types of self-disclosures they contain. We attempted to assign both theory-based and automatic category assignments to each comment in our dataset to represent the type of information disclosed by annotators. Our theory-based assignment was informed in part by our intuition and related psychological theory, and in part by our exploratory analysis of the automatic assignments.
@@ 3.1 Automatic Categorization
We started with both LDA topics Blei et al. (2003) and k-means clusters as ways of grouping comments. When clustering and topic modeling, we use only posts that contain a self-disclosure from the list in Appendix A so that clusters relate more to personal statements. This leaves us with 570k posts, or 16% of the original data. About 400 authors have never written a post with a self-disclosure phrase and were excluded from our category experiments. Clusters were created using post embeddings generated using SBERT with the same parameters discussed in §4. Both approaches provide a method by which a single group label (topic or cluster) can be assigned to each comment. We found through manual inspection in preliminary experiments that both provided reasonable groupings and decided to proceed with k-means, as it is more easily interpretable. We also experimented with changing the number of clusters by multiples of five up to 25, finding that ten provided a reasonably interpretable and coherent set of clusters.
@@
Table 4: Situation split performance (Accuracy and F1) across varying numbers of maximum samples, with best performance in bold. Similar 5 comments to random 5 comments statistically significant (permutation test, $p<0.0002$).
@@ 3.2 Theory-based Categorization
Building upon the Social Penetration Theory for self disclosure Altman (1973), we introduce a theory-based categorization framework for understanding the types of self-disclosures that occur during interpersonal communication. Our taxonomy of annotator self-disclosure statements includes (a) Demographics, which was found to be the most common in other studies on annotator modeling, (b) Experiences, which includes generally self-descriptive statements about one’s past experiences and relatively more stable traits, (c) Attitudes, which are self-descriptive statements of social views, morals, values and belief systems, and (d) Relationships, which describe the sociability and familial relations of the annotator. The clustering methods in all preliminary experiments picked up on the relationship between those in conflict, similar to what was found in Welch et al. (2022). We analyzed the most common n-grams around self-disclosure phrases, grouped authors by post count, tagged parts of speech to identify frequent patterns and verbs, and used Stanza’s coreference resolution (Qi et al., 2020) to clarify sentence meaning and better understand the main themes in self-disclosure. Theory-based and automatic approaches group self-disclosures by the type of information and the context (e.g. affect or conflict type) in which the disclosure was made, respectively. Post and author counts for each category are listed in Tables 2 and 3. We extended the age and gender regular expressions from Welch et al. (2020) and Plepi et al. (2022) to include a variety of ways of expressing age, non-binary and transgender, along with age and gender combinations, e.g. (24F representing 24 year old female). Our set of initial self-disclosure statements was expanded from Plepi et al. (2024) and similar to Mazaré et al. (2018), but without enforcing part-of-speech constraints. We also borrowed expressions from Tigunova et al. (2020), which included age, gender, family status, profession, and hobby, but expanded the set of cases. The full set of regular expressions we developed is listed in Table 6 in the Appendix.
@@
After using regex to extract the self-disclosure information, we randomly selected 100 comments from each group to annotate by hand. We found that there were 9 comments in the demographic group, 17 posts in the attitudes group, 25 in the experiences group, and 11 in the relationships group which were not clean self-disclosure phrases. For example, the attitudes regex patterns currently capture the phrases: “I think driving is scary” and “I think thats a good plan”. The first phrase gives a clear opinion and is fully contained. The second sentence gives an opinion on something that has been mentioned outside of the current comment. We were initially expecting to extract clear opinions with this pattern, but in reality self-contained disclosures are difficult to extract. Note that the use of an LLM could potentially help, but would introduce another source of error, while requiring substantially more compute effort for a large corpus. Future work could likely improve upon our findings by improving the extraction approach for self-disclosures, possibly through a multi-stage filtering process.
@@ 4 Methodology
As a baseline, we reproduced the results of Plepi et al. (2022) for their situation split using the same parameters from their setup. They showed that the situation split was the most challenging, i.e. predicting how annotators respond to new conflict posts (see Appendix C for details on other splits). We used DistilRoBERTa Sanh et al. (2019) for the base SBERT classifier Reimers and Gurevych (2019) with a hidden size 768 and max length 512. We use an added linear layer for classification, trained the model for 10 epochs with the Adam optimizer, learning rate 1e31e-3, and focal loss Lee et al. (2021).
Experimental Setup. The brief exploration of self-disclosure statements in previous work on generation raised several questions for us. RQ1: Which sampling parameters perform best? First, how many samples are ideal for classifying judgments? Second, are similar samples more useful than random samples? Lastly, is selecting sentences more effective than taking the whole posts? We hypothesized that comments similar to the post we are classifying should perform better. Given the results of Plepi et al. (2024), we assumed that 15-20 similar posts would work best. However, we hypothesized that individual sentences may work better because they focus on only the important information and eliminate potential noise in longer posts.
@@
Our next question was whether or not manually categorizing personal information was necessary, or if we could rely on automatic methods. RQ3: Does automatically categorizing self-disclosures via clustering perform as well as theory-based categorization? We hypothesized that clusters would underperform theory-based categories because clustering makes it difficult to enforce that similarity among clustered comments is driven mainly by author related information. rather than by topical or surface level features. Furthermore, we hypothesized that the best representation of an individual should include a diverse array of information about that person. If we sample from different manual and automatic categories, we expect performance to increase.
To perform the first set of experiments we used all comments from all annotators, meaning that some randomly sampled or similar comments/sentences may not belong to any theory-based or automatic category. We chose to investigate the sampling parameters first (RQ1), as the search space grows exponentially if we combine this with RQ2 and RQ3. For the latter research questions, we sampled only similar posts from one category at a time. For each category, we use the same set of annotators, however, each annotator does not have the same number of comments in each category, so we used up to five.
All of the models in our experiments were trained using a single NVIDIA H100 GPU. Generating similar comment embeddings and sentence tests took approximately 5 hours and only needed to be computed once each. Generating the embeddings for each category took less than a minute. A new model was needed for each of our experimental settings, a total of 24 models for the experiments in Table 4, a total of 14 models for the experiments in Table 5. Each model took approximately 2 hours to train and the results were averaged over 5 runs.
@@
Diverse Sampling
Both Categorizations - 68.1 56.0
Theory Based Categorization - 67.6 56.7
Automatic Categorization - 66.3 55.6
Table 5: Accuracy and F1 scores for annotator models trained on the baseline, theory based, automatic, and diverse grouping methods. 5+% indicates the percentage of annotators who had at least 5 posts for the category. All comments baseline represents previous work by Plepi et al. (2022). Bold numbers are statistically significant in comparison to the no comments baseline (permutation test, $p<0.05$), highest numbers per subsection are underlined.
@@
5 Results
The results of our experiments to determine the optimal strategies and parameters for sampling information about annotators (RQ1) are displayed in Table 4, showing accuracy and macro F1 score. We find that training on sentences underperforms using the full comments. It could be that after a certain point, adding more sentences or comments stops adding useful information. When the texts are sorted by similarity, there is a trade-off, where adding more similar text results in adding less relevant information. When increasing the number of similar sentences, we are adding a small amount of similar information, however the results from the model using 20 similar sentences already show a drop in performance because each sentence is less similar and less useful than the last. Overall, we find that adding the 5 most similar comments performs best. Using similar information, not surprisingly, is more effective than adding random posts or sentences. Using 5 similar comments significantly outperforms a baseline which uses all posts and which uses no posts (t-test, $p<0.0001$). We use the setting of 5 most similar comments for subsequent experiments in Table 5.
After identifying the optimal sampling parameters, we proceeded to address RQ2 and RQ3. Our results for experiments with different categories are shown in Table 5, with highest accuracy and F1 in bold separately for theory-based, automatic, and diverse categories. The highest performing theory-based category was demographics, followed by attitudes, and finally experiences, and relationships. We compared each result to a baseline that uses no comments from the annotators and a baseline that uses all comments, as in previous work. Unsurprisingly, no comments performs worst, while all comments achieves stronger performance while using substantially more data. The highest performing automatic clusters were C7: Sympathy & Support and C9: Negative Affect. C3: Parenting & Discipline had a notably high accuracy. Note that some of the higher performing categories were not significantly different. Notably, C10 had one of the lowest coverage rates with only 23% of authors having five or more posts in that cluster. Diverse sampling performed competitively, with theory-based categorization outperforming automatic clusters or a combination of the two.
33.6020204040606080801001001.20.50.5111.51.522
050100AgeGenderIdentity050100HobbyPossessionWork